OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (kagıtucak):

Divide both sides with -3. |x+7|=4 x+7=4 x=-3 and x+7=-4 x=-11 {-3,-11}

OpenStudy (anonymous):

thanks

OpenStudy (kagıtucak):

nop